# Who Should I Listen To? Adaptive Collaboration in Personalized Federated Learning

### v2

@@ Abstract
Data heterogeneity is a central challenge in federated learning, and personalized federated learning (PFL) aims to address it by tailoring models to each client’s distribution. Yet many PFL methods fail to consistently outperform local or centralized baselines, suggesting a mismatch between the collaboration they enforce and the structure of the data. We propose an approach based on adaptive collaboration, where clients decide adaptively not only how much to rely on others, but also whom to trust—at the level of individual examples. We instantiate this principle in FedMosaic, a federated co-training method in which clients exchange predictions over a shared unlabeled dataset. This enables fine-grained trust decisions that are difficult to achieve with parameter sharing alone. Each client adjusts its loss weighting based on the agreement between private and public data, and contributes to global pseudo-labels in proportion to its estimated per-example confidence. Empirically, FedMosaic improves upon state-of-the-art PFL methods across diverse non-IID settings, and we provide convergence guarantees under standard assumptions. Our results demonstrate the potential of data-aware collaboration for robust and effective personalization.
@@ 1 Introduction
Table 1: Average test Accuracy on DomainNet and Office dataset (details in sec.4). Most personalized FL methods fail to surpass local training baseline. FedMosaic consistently exceeds both core baselines through adaptive, example-level collaboration. Color Map: baselines, worse than baselines, worse than local training, better than baselines.
@@
Centralized $0.6624$ $(0.004)$ $0.4092$ $(0.006)$
FL FedAvg $0.3100$ $(0.008)$ $0.3725$ $(0.008)$
FedProx $0.5523$ $(0.001)$ $0.5839$ $(0.003)$
PFL Per-FedAvg $0.7248$ $(0.004)$ $0.7192$ $(0.005)$
pFedMe $0.7521$ $(0.005)$ $0.7483$ $(0.007)$
APFL $0.8059$ $(0.003)$ $0.8091$ $(0.001)$
FedPHP $0.7825$ $(0.006)$ $0.7636$ $(0.004)$
Local Training $0.8464$ $(0.001)$ $0.8679$ $(0.004)$
FedMosaic $\mathbf{0.8744}$ $(0.0002)$ $\mathbf{0.8906}$ $(0.0001)$
Federated learning promises collaboration without compromising privacy, but when client data is heterogeneous, standard approaches often collapse. Personalized federated learning (PFL) seeks to address this challenge, yet much of the field rests on shaky foundations. While the task of PFL is well-defined in principle, i.e., optimizing performance on each client’s distribution, many existing methods are evaluated under ill-posed experimental conditions. Several fail to compare against purely local training (Fallah et al., 2020; T Dinh et al., 2020; Deng et al., 2020; Li et al., 2021d), where clients learn independently without communication. If a PFL method underperforms this baseline, collaboration is not only unnecessary but actively detrimental (cf. Tab. 1). Similarly, many approaches omit comparisons to centralized training, which, while often infeasible due to privacy constraints, reveals whether a global model could succeed despite heterogeneity. In cases where centralized training performs well, the real problem lies not in the need for personalization, but in the failure of existing federated aggregation schemes under distribution shift.
In personalized federated learning, each client $i\in[N]$ has access to a dataset $\mathcal{D}_{i}$ drawn from a local data distribution $\mathcal{D}_{i}$, which may vary across clients. The goal is to learn personalized models $h_{1},\dots,h_{N}$ that minimize the local risk at each client, that is,
$\min_{h_{1},\dots,h_{N}}\sum_{i=1}^{N}\mathbb{E}_{(x,y)\sim\mathcal{D}_{i}}\left[\mathcal{L}\left(h(x),y\right)\right]\enspace.$
This objective highlights a subtle but crucial point: since clients are evaluated on their own distributions, local training can sometimes outperform centralized training. Thus, strong local baselines must be considered when assessing the true benefit of personalization.
One might object that centralized models optimize global performance, while PFL focuses on client-specific distributions. However, this overlooks a key tradeoff. Centralized models benefit from far more data, which can reduce variance and improve generalization, even on heterogeneous test sets. Personalized models may better align with individual distributions, but often suffer from data scarcity and overfitting. In practice, centralized training often outperforms PFL methods when the gains from data volume outweigh the benefits of specialization. For this reason, centralized training remains a critical reference point, even when personalization is the objective.
@@
This confusion arises because personalization is often proposed as a vague remedy for heterogeneity without clarifying when it is needed, how it should work, or what the right metrics for success are. We argue that progress in PFL requires a clearer conceptual foundation. Personalization should not be the default; it should emerge from an understanding of what each client needs and how collaboration benefits them. In particular, a meaningful PFL solution must adapt the degree of collaboration based on the client’s data and context, rather than enforcing a uniform global training scheme. It must also account for heterogeneity not only at the client level but at the level of individual examples or regions of the data space, since clients may align on some concepts and diverge on others. Without such granularity, personalization reduces to blunt model-level adjustments that fail to capture the true complexity of data heterogeneity.
These considerations point to a general principle: effective personalization in federated learning must enable adaptive, data-specific collaboration. A PFL method should allow clients to modulate how much they learn from others based on their own needs, while supporting fine-grained interactions that reflect the structure of the data itself. While this principle could be realized in many ways, one particularly promising direction comes from federated semi-supervised learning, e.g., distributed distillation (Bistritz et al., 2020) or federated co-training (Abourayya et al., 2025). Unlike classical methods that share model parameters, these methods enable clients to interact by sharing predictions on a shared, unlabeled public dataset. This shift makes it possible to personalize collaboration per example: a client can contribute only to those examples for which it has expertise, and consensus mechanisms can weight contributions accordingly. Such mechanisms naturally reflect the varying reliability of clients across the data space. Two clients who have seen many cat images might agree on how to label a new cat photo, while a client that has never seen dogs should not influence the labeling of dog images. This level of collaboration is fundamentally difficult—if not impossible—to achieve through parameter aggregation alone.
@@
In this work, we demonstrate that this principle can be put into practice. We propose a method called FedMosaic, a personalized federated co-training algorithm that enables adaptive, fine-grained collaboration through prediction sharing. FedMosaic uses two key mechanisms: a dynamic weighting strategy that allows each client to adjust its reliance on global versus local signals, and an expertise-aware consensus mechanism that weights client contributions according to their demonstrated competence on different parts of the data. Both mechanisms operate on predictions over a public dataset, enabling personalization that is responsive to real distributional structure rather than imposed heuristically.
While FedMosaic achieves strong empirical performance, outperforming state-of-the-art personalized FL methods as well as both local and centralized baselines, its core contribution is conceptual. It shows that meaningful personalization is achievable when we rethink the way clients interact. Rather than pushing more complex model-sharing schemes or training tricks, we argue for a shift in perspective: personalization is not a layer added on top of existing FL algorithms, but a question of how collaboration happens. FedMosaic is one instantiation of this broader view, illustrating that principled, adaptive collaboration can substantially advance personalized federated learning. We hope this perspective inspires future research on data-driven personalization and more nuanced models of federated interaction.
@@ 2 Related Work
Personalized Federated learning (PFL)
@@
Meta-learning and Regularization-Based Methods: The methods optimize a shared initialization or constrain local updates. For example, Per-FedAvg (Fallah et al., 2020)uses MAML to learn a shared initialization, while Ditto (Li et al., 2021b) regularizes local updates toward a global reference. PFedMe(T Dinh et al., 2020) applies bi-level optimization to decouple personalization from global learning.
Personalized Aggregation strategies: These methods dynamically aggregate models based on client similarity or adaptive weighting. APFL (Deng et al., 2020) introduces an adaptive mixture of global and local models, allowing clients to interpolate between shared and personalized parameters based on their data distribution. FedAMP(Huang et al., 2021) uses attention to weight client contributions based on similarity. Other methods select collaborators (e.g., FedFomo (Zhang et al., ), FedPHP (Li et al., 2021d)) or apply layer-wise attention (FedALA (Zhang et al., 2023c)).
Model Splitting Architectures: These approaches partition models into shared and personalized components. FedPer (Arivazhagan et al., 2019)eeps shared base layers and personalizes top layers. FedRep shares a backbone but personalizes the head.(Collins et al., 2021) shares a backbone but personalizes the head. FedBN (Li et al., 2021c) personalizes batch normalization layers to tackle feature shift. Other recent methods such as FedAS (Yang et al., 2024), GPFL (Zhang et al., 2023b), and FedBABU (Oh et al., 2021) disentangle or freeze specific parts of the model to balance generalization and personalization. PFedHN (Shamsian et al., 2021) uses a hypernetwork that generates personalized model parameters conditioned on client identity.
Knowledge Distillation Approaches: Knowledge distillation methods transfer knowledge from global or peer models to personalized local models. FedProto (Tan et al., 2022) aligns class-wise feature prototypes across clients, FedPAC (Xu et al., 2023) uses contrastive learning to distill knowledge into personalized models, and FedKD (Wu et al., 2022) reduces communication cost by distilling knowledge from a teacher ensemble to lightweight client models. FedMatch (Chen et al., 2021) uses consistency regularization to unlabeled and noisy data, FedDF (Lin et al., 2020) aggregates predictions via ensemble distillation, and FedNoisy (Liang et al., 2023) focuses on robust aggregation in the presence of noisy labels or adversarial participants.
@@
Despite this progress, existing PFL methods often share several limitations: (i) Static collaboration: Most PFL methods rely on fixed rules (e.g., aggregation weights or model splits), lacking adaptivity to client-specific or example-level variation. (ii) Privacy risks: Sharing model parameters, gradients, or even soft labels may expose sensitive information. (iii) Limited generality: Many methods are tailored to specific heterogeneity types (e.g., label skew in case of FedMix, or feature shift in case of FedBN). (iv) Communication/computational overhead: Some require complex multi-model training or costly synchronization. To overcome these limitations, we argue that PFL methods should use some form of dynamic modulation and per-example trust weighting. We propose a method that achieves both while sharing only hard predictions on an unlabeled dataset to enhance privacy and minimize overhead.
@@ 3 Personalized Federated Co-Training: Adaptive and Expert-Aware Collaboration
We now introduce FedMosaic as a concrete realization of the principle proposed in Sec. 1, i.e., that effective personalization in federated learning arises from adaptive, data-specific collaboration.
It builds on the idea of federated co-training, a privacy-friendly alternative to parameter sharing. In federated co-training (Abourayya et al., 2025), each client shares only its predictions (hard labels) on a shared, unlabeled public dataset. These predictions are aggregated into a consensus pseudo-labeling, which is then used to augment the local datasets. This approach has been shown to substantially improve privacy. Therefore, in our setting we assume a shared unlabeled dataset $U\subset\mathcal{X}\times\mathcal{Y}$. Such datasets are often available or can be synthesized, for example from public data repositories or via generative models. Clients share only their predictions on $U$ and no sensitive information or parameters are exchanged.
However, this basic setup introduces two key challenges. (i) When to trust the global signal? Clients may disagree with the global consensus, particularly when their data distribution diverges from the majority. Blindly following the global pseudo-labels can harm performance. (ii) Whose predictions to trust? Not all clients are equally competent across all regions of the data space. A good method should rely more on clients that are confident and accurate for specific examples.
To address these challenges, FedMosaic introduces two mechanisms, dynamic loss weighting and confidence-based aggregation.
In dynamic loss weighting, each client learns how much to trust the global signal versus its own local predictions. This tradeoff is modulated via an adaptive weight. Then, client $i$ in round $t$ minimizes the objective.
$\mathcal{L}_{i}^{t}(h)=\mathcal{L}(h,D_{i})+\lambda_{i}^{t}\cdot\mathcal{L}(h,P_{t})$
for a model $h$, where $\lambda_{i}^{t}$ is the dynamic weight and $P_{t}=(U,L_{t})$ is the pseudo-labeled public dataset in round $t$. We want $\lambda$ to be $1$ when the model performs equally well on local and global data to ensure collaboration, and to be 00 when the model performs poorly on global data, indicating that collaboration should be avoided. In cases where local data is a noisy variant of global data, $\lambda$ might even exceed $1$, but it should be bounded. To fulfill these criteria, we propose to define it as
$\lambda_{i}^{t}=\exp\left(-\frac{\mathcal{L}(h^{i}_{t-1},P_{t})-\mathcal{L}(h^{i}_{t-1},D_{i})}{\mathcal{L}(h^{i}_{t-1},D_{i})}\right)\in[0,e].$
The ratio measures the relative difference in performance on local and global data and the exponential transformation ensures that $\lambda$ is positive, smooth and bounded. This allows clients to balance collaboration and independence based on their own performance.
In confidence-based aggregation, the predictions of each client are weighed by their per-sample confidence, or expertise (in contrast to treating all clients equally when generating pseudo-labels). Clients contribute more to the global consensus when they are likely to be correct, and less when uncertain. This encourages accurate clients to lead the collaboration in regions where they are competent, while reducing the impact of noisy or misaligned peers. In practice, this means that every $b$ rounds, each client not only shares its predictions $L_{t}^{i}\in\{0,1\}^{|U|\times C}$ on the publica dataset $U$ as a one-hot encoded matrix, but also an associated confidence vector $E_{t}^{i}\in(0,\infty)^{|U|}$, where $E_{t}^{i}[j]$ quantifies the estimated reliability of client $i$’s prediction on $x_{j}$. The server collects all $(L_{t}^{i},E_{t}^{i})$ and computes a weighted score matrix
$S_{t}=\sum_{i=1}^{m}\operatorname{diag}(E_{t}^{i})\cdot L_{t}^{i}\in\mathbb{R}^{|U|\times C}.$
With this score matrix, the consensus is now formed via the confidence-based aggregation
$L_{t}[j]=\arg\max_{c\in[C]}S_{t}[j,c],\quad\forall j\in\{1,\dots,|U|\}\enspace.$
These consensus pseudo-labels are then redistributed to all clients for further training.
We propose two variants of confidence scores $E_{t}^{i}[j]$, namely a class-frequency-based heuristic, where $E_{t}^{i}[j]$ for example $x_{j}$ predicted to belong to class $y$ is the frequency of class $y$ in the local training set of client $i$, and (2) an uncertainty-based score, where $E_{t}^{i}[j]$ is inversely related to the predictive entropy or margin of $h_{t}^{i}(x_{j})$, i.e., the higher the score, the more certain the client about its prediction. Both choices allow the aggregation to selectively trust clients that are more reliable on specific regions of the data space. We evaluate both variants in Section 4. The full algorithm is provided in Algorithm 1, where differences to the original federated co-training are highlighted in green.
@@
1 initialize local models $h_{0}^{1},\dots,h_{0}^{m}$,  $P\leftarrow\emptyset$
2
3Locally at client $i$ at time $t$ do

### v3

# Whom to Trust? Adaptive Collaboration in Personalized Federated Learning

@@ Abstract
Data heterogeneity poses a fundamental challenge in federated learning (FL), especially when clients differ not only in distribution but also in the reliability of their predictions across individual examples. While personalized FL (PFL) aims to address this, we observe that many PFL methods fail to outperform two necessary baselines, local training and centralized training. This suggests that meaningful personalization only emerges in a narrow regime, where global models are insufficient, but collaboration across clients still holds value. Our empirical findings point to two key ingredients for success in this regime: adaptivity in collaboration and fine-grained trust, at the level of individual examples. We show that these properties can be achieved within federated semi-supervised learning, where clients exchange predictions over a shared unlabeled dataset. This enables each client to align with public consensus when it is helpful, and disregard it when it is not, without sharing model parameters or raw data. As a concrete realization of this idea, we develop FedMosaic, a personalized co-training method where clients reweight their loss and their contribution to pseudo-labels based on per-example agreement and confidence. FedMosaic outperforms strong FL and PFL baselines across a range of non-IID settings, and we prove convergence under standard smoothness, bounded-variance, and drift assumptions. In contrast to many of these baselines, it also outperforms local and centralized training. These results clarify when federated personalization can be effective, and how fine-grained, trust-aware collaboration enables it.
@@ 1 Introduction
Federated learning (FL) enables collaborative machine learning across distributed data sources without direct data sharing. Classical methods such as FedAvg (McMahan et al., 2017), aim to train a single global model across all clients. This approach can succeed when data distributions are sufficiently similar, but collapses under strong distributional shifts. In highly heterogeneous settings, the promise of collaboration breaks down: models trained jointly may perform worse than models trained independently.
Personalized Federated Learning (PFL) addresses this challenge by shifting the goal. Rather than optimizing a shared global model, the goal is to use collaboration to improve each client’s personalized model. For example, Tab. 1 shows that in heterogeneous regimes both FL and even centralized training perform worse than local training,i.e., clients learning independently without any communication. This underlines the requirement for PFL, but also highlights an often-overlooked baseline: when no method outperforms local training, collaboration is not just ineffective—it is detrimental. Yet many PFL methods fail to beat this baseline (cf. Tab. 1), casting doubt on their utility.
Table 1: Average test Accuracy on DomainNet and Office-10 dataset (details in sec.4). Most personalized FL methods fail to surpass local training baseline. FedMosaic exceeds both core baselines through adaptive, example-level collaboration. Color Map: baselines, worse than baselines, worse than local training, better than baselines.
@@
Centralized $66.24$ $(0.4)$ $40.92$ $(0.6)$
FL
FedAvg $31.00$ $(0.8)$ $37.25$ $(0.8)$
FedProx $55.23$ $(0.1)$ $58.39$ $(0.3)$
PFL
Per-FedAvg $72.48$ $(0.4)$ $71.92$ $(0.5)$
pFedMe $75.21$ $(0.5)$ $74.83$ $(0.7)$
APFL $80.59$ $(0.3)$ $80.91$ $(0.1)$
FedPHP $78.25$ $(0.6)$ $76.36$ $(0.4)$
Local Training $84.64$ $(0.1)$ $86.79$ $(0.4)$
FedMosaic $\mathbf{87.44}$ $(0.02)$ $\mathbf{89.06}$ $(0.01)$
@@
This widespread failure to measure true collaborative gain arises because ”personalization” is often treated as a vague remedy for heterogeneity without a clear underlying principle. We argue that progress requires a new foundation. Personalization shouldn’t be a default modification to an existing FL algorithm; it should emerge from a principled understanding of what each client needs and how collaboration can help. A meaningful PFL solution must adapt the degree and nature of collaboration based on client context. It must also account for heterogeneity not just between clients, but at the level of individual examples. Clients may align on some concepts (e.g., identifying cats) and diverge on others (e.g., identifying specific dog breeds), and collaboration should reflect this granularity.
In formal terms, PFL aims to minimize the sum of local risks across $m$ clients with heterogeneous data distribution $\mathcal{D}_{i}$ and personalized models $h_{1},\dots,h_{m}$:
$\min_{h_{1},\ldots,h_{m}}\sum_{i=1}^{m}\mathbb{E}_{(x,y)\sim\mathcal{D}_{i}}\left[\mathcal{L}\left(h_{i}(x),y\right)\right]\kern 5.0pt.$
In this setting, local model may outperform global or centralized models, making strong local and centralized baselines éssential. The key trade-off between the massive data access of a centralized model versus the specialization of a local one, is the central tension PFL must navigate in an adaptive and data-specific way.
While federated learning can adapt by weighing parameters according to similarity (Huang et al., 2021; Zhang et al., 2021), data-specific collaborations require a shift in mechanism. Rather than aggregating model parameters, we propose to use federated semi-supervised learning (Bistritz et al., 2020; Abourayya et al., 2025) where clients share predictions on a public dataset. Collaboration is achieved by enforcing consensus between clients. We propose to adapt this consensus mechanism so that clients can contribute only on examples where they have expertise and can selectively trust others based on their demonstrated competence. Two clients familiar with cats can confidently collaborate on a new cat photo, while a client that has only seen cars should not influence the labeling of cat images. This form of selective, example-level trust is fundamentally difficult to achieve through parameter averaging alone.
@@
In this work, we demonstrate this principle in practice. We propose a personalized Federated Co-Training approach (FedMosaic) that enables adaptive, fine-grained collaboration through two key mechanisms: a dynamic weighting strategy allowing clients to balance global and local signals in each communication round, and an expertise-aware consensus mechanism that weights peer contributions by their competence on different data regions. Both mechanisms operate on predictions over a public dataset, enabling personalization that is responsive to the data’s true structure.
While FedMosaic achieves state-of-the-art empirical performance across benchmarks, its main contribution is conceptual. It redefines personalization as a question of collaborative structure, not just algorithm design. Our results show that principled, example-level collaboration can unlock the full potential of personalized federated learning.
@@ 2 Related Work
Personalized Federated learning (PFL):
@@
Meta-learning and Regularization-Based Methods optimize a shared initialization or constrain local updates. E.g., Per-FedAvg (Fallah et al., 2020) learns a shared initialization, while Ditto (Li et al., 2021b) regularizes local updates toward a global reference. PFedMe (T Dinh et al., 2020) applies bi-level optimization to decouple personalization from global learning. Personalized Aggregation strategies dynamically aggregate models based on client similarity or adaptive weighting. APFL (Deng et al., 2020) introduces an adaptive mixture of global and local models, allowing clients to interpolate between shared and personalized parameters based on their data distribution. FedAMP (Huang et al., 2021) uses attention to weight client contributions based on similarity. Other methods select collaborators (e.g., FedFomo (Zhang et al., 2021), FedPHP (Li et al., 2021d)) or apply layer-wise attention (FedALA (Zhang et al., 2023c)). Model Splitting Architectures partition models into shared and personalized components. FedPer (Arivazhagan et al., 2019)eeps shared base layers and personalizes top layers. FedRep shares a backbone but personalizes the head.(Collins et al., 2021) shares a backbone but personalizes the head. FedBN (Li et al., 2021c) personalizes batch normalization layers to tackle feature shift. Other recent methods such as FedAS (Yang et al., 2024), GPFL (Zhang et al., 2023b), and FedBABU (Oh et al., 2021) disentangle or freeze specific parts of the model to balance generalization and personalization. PFedHN (Shamsian et al., 2021) uses a hypernetwork that generates personalized model parameters conditioned on client identity. Knowledge Distillation Approaches transfer knowledge from global or peer models to personalized local models. FedProto (Tan et al., 2022) aligns class-wise feature prototypes across clients, FedPAC (Xu et al., 2023) uses contrastive learning to distill knowledge into personalized models, and FedKD (Wu et al., 2022) reduces communication cost by distilling knowledge from a teacher ensemble to lightweight client models. FedMatch (Chen et al., 2021) uses consistency regularization to unlabeled and noisy data, FedDF (Lin et al., 2020) aggregates predictions via ensemble distillation, and FedNoisy (Liang et al., 2023) focuses on robust aggregation in the presence of noisy labels or adversarial participants. PerFed-CKT (Cho et al., 2021) enhances personalization by clustering clients with similar data distributions and facilitating knowledge transfer through logits instead of model parameters. Jeong & Kountouris (2023) proposes a fully decentralized PFL framework where clients share distilled knowledge with neighboring clients, enabling personalization without a central server.FedD2S (Atapour et al., 2024) introduces a data-free federated knowledge distillation approach that employs a deep-to-shallow layer-dropping mechanism.
@@
Despite this progress, existing PFL methods often share several limitations: (i) Static collaboration: Most PFL methods rely on fixed rules (e.g., aggregation weights or model splits), lacking adaptivity to client-specific or example-level variation. (ii) Privacy risks: Sharing model parameters, gradients, or even soft labels may expose sensitive information. (iii) Limited generality: Many methods are tailored to specific heterogeneity types (e.g., label skew in case of FedMix, or feature shift in case of FedBN). (iv) Communication / computational overhead: Some require complex multi-model training or costly synchronization. To overcome these limitations, we argue that PFL methods should use some form of dynamic modulation and per-example trust weighting.
@@ 3 Personalized Federated Co-Training: Adaptive and Expert-Aware Collaboration
We now introduce Personalized Federated Co-Training (FedMosaic), a concrete realization of the principle that effective personalization arises from adaptive, data-specific collaboration. Our method builds upon the framework of federated co-training (Abourayya et al., 2025), a privacy-preserving paradigm where clients collaborate by sharing hard predictions on a shared, unlabeled public dataset, $U$ (we analyze the impact of this dataset’s size and distribution in sec.4). This process creates a consensus pseudo-labeled dataset, which clients use to augment their local training.
While this approach avoids sharing sensitive model parameters and soft labels, it introduces two critical challenges for personalization:
  1. 1.
    When to trust the global signal? A client’s local data may conflict with the global consensus. Blindly trusting pseudo-labels can harm a model that is already well-specialized.
  2. 2.
    Whose predictions to trust? Clients possess varying levels of expertise across the data space. A naive consensus that treats all clients equally will be corrupted by noisy or misaligned predictions.
FedMosaic addresses these challenges directly with two core mechanisms: (1) dynamic loss weighting, which allows each client to adaptively decide when to trust the global signal, and (2) confidence-based aggregation, which intelligently decides whose predictions to trust.
Dynamic Loss Weighting: Deciding When to Trust: To allow clients to autonomously balance global collaboration with local specialization, we introduce a dynamic weight $\lambda_{i}^{t}$, into the local objective. At each round $t$, client $i$ minimizes the combined loss:
$\mathcal{L}_{i}^{t}(h)=\mathcal{L}\left(h,D_{i}\right)+\lambda_{i}^{t}\cdot\mathcal{L}\left(h,P_{t}\right)$
where $\mathcal{D}_{i}$ is the client’s private data and $P_{t}$ is the pseudo-labeled public dataset.The weight $\lambda_{i}^{t}$ modulates the influence of the global signal.Our choice of the function for computing $\lambda_{i}^{t}$was driven by the need for a smooth, bounded, and interpretable mechanism. We define it as:
$\lambda_{i}^{t}=\exp\left(-\frac{\mathcal{L}\left(h_{t-1}^{i},P_{t}\right)-\mathcal{L}\left(h_{t-1}^{i},D_{i}\right)}{\mathcal{L}\left(h_{t-1}^{i},D_{i}\right)}\right)$
This exponential form satisfies several desirable properties. It ensures positivity $\left(\lambda_{i}^{t}>0\right),$ avoids discontinuities, and smoothly adjusts the client’s trust based on the relative performance of its model on global versus local data. The behavior is highly intuitive:
  • Conflict $\left(\mathcal{L}_{\text{global }}\gg\mathcal{L}_{\text{local }}\right)$: If the consensus pseudo-labels are harmful, the global loss term increases, causing $\lambda_{i}^{t}\rightarrow 0$ and prompting the client to rely on its local data.
  • Alignment $\left(\mathcal{L}_{\text{global }}\approx\mathcal{L}_{\text{local }}\right)$ : If the consensus is helpful and aligns with local data, $\lambda_{i}^{t}\approx 1$ achieving a balance between personalization and collaboration.
  • Enhancement $\left(\mathcal{L}_{\text{global }}<\mathcal{L}_{\text{local }}\right)$: If the consensus provides a cleaner signal than the noisy local data, $\lambda_{i}^{t}>1$, encouraging the client to trust the collaborative signal more heavily.
Confidence-Based Aggregation: Deciding Whose to Trust: To address the varying expertise of clients, we replace the standard uniform aggregation of predictions with a confidence-based consensus. Instead of just sharing hard labels, each client $i$ also communicates a confidence vector $E_{t}^{i}\in(0,\infty)^{|U|}$, where $E_{t}^{i}[j]$ quantifies its estimated expertise on its prediction for example $x_{j}\in U$.The server then computes a weighted score matrix $S_{t}$ by aggregating the one-hot predictions $L_{t}^{i}$ from each client, weighted by their corresponding expertise:
$S_{t}=\sum_{i=1}^{m}\operatorname{diag}\left(E_{t}^{i}\right)\cdot L_{t}^{i}\in\mathbb{R}^{|U|\times C}$
The final consensus pseudo-label for each example is determined by the highest aggregate score:
$L_{t}[j]=\arg\max_{c\in[C]}S_{t}[j,c],\quad\forall j\in\{1,\ldots,|U|\}$
This mechanism allows clients who are more confident or reliable about specific data regions to have a greater influence on the consensus, effectively reducing the impact of noise from non-expert clients. We explore two practical instantiations for the confidence scores $E_{t}^{i}$: a class-frequency-based heuristic and an uncertainty-based score derived from the model’s predictive entropy. The full procedure is detailed in Algorithm  1.
@@
1
20.1cm initialize local models $h_{0}^{1},\dots,h_{0}^{m}$,  $P\leftarrow\emptyset$
3
4Locally at client $i$ at time $t$ do